a 3edl OS sH Nonpolar auy paeg H H и jo qun Polyatomic ion IWAVANAL. HI-POLYMER Instructor Sectid. Nat essacs ERASER
Answer #1

CHY 7 • Number of valence electron un Chy is & • Lewis e Valence electron sphere of CH4. centructure neocon - Molecular Geome180 Molecular Geometry no of lone pour on Carbono no. of Sigma bond pair = 2 Tolal e bain = 0+1 = 2 (3p) Molecular Geometry i→ The electronegativity difference betwen both atoms is less than 0.4. Hence bonds are non polar . If bonds are non-polar , w